# TritonRL: Training LLMs to Think and Code Triton Without Cheating

### v1

@@ F.2 Example of invalid Triton codes in KernelBook
F.3 Example of invalid Triton codes for Level 2 tasks
The following is an example of invalid Triton code generated during RL training of TritonRL. This level 2 task aims to implement a two-branch linear transformation with message passing, structured like a Graph Neural Network (GNN) layer. However, while the code implment linear transformations using a custom Triton kernel, it fails to fully realize the intended GNN functionality with only Triton kernels by using torch.bmm.
@@ F.3 Example of invalid Triton codes for Level 2 tasks
F.4 Invalid Triton code examples
@@
channel_idx = (offsets // stride_channel) % out_C
@@
F.5 SFT and RL Dataset construction with KernelBook
@@
For each given PyTorch reference code, we construct an instruction for DeepSeek-R1 to generate CoTs and Triton kernels as:
@@
Given the instruction for each PyTorch reference code, we collect (CoT, Triton kernel code) pairs fom DeepSeek-R1 and construct outputs for SFT by concatenating the pairs as follows:
@@
In this manner, for each Pytorch reference code in KernelBook, we construct 5 (input, output) SFT samples.
@@
Appendix G Evaluation and examples
G.1 Evaluation with KernelBench
@@
G.2 Example of Triton codes with speedup >1>1
@@
Problem 61 in KernelBench Level 1: 3D Transposed Convolution with x1.5 speedup.
@@
def conv_transpose3d_kernel(
input_ptr, weight_ptr, output_ptr,
B, IC, OC, D_in, H_in, W_in, D_out, H_out, W_out,
input_batch_stride, input_channel_stride, input_d_stride, input_h_stride, input_w_stride,
weight_in_channels_stride, weight_out_channels_stride, weight_d_stride, weight_h_stride, weight_w_stride,
output_batch_stride, output_channel_stride, output_d_stride, output_h_stride, output_w_stride,
@@
# Each block handles BLOCK_SIZE output elements
pid = tl.program_id(0)
n_elements = B * OC * D_out * H_out * W_out
@@
# Mask to avoid out-of-bounds access in last block
@@
# Precompute for flattening and indexing
OC_D_outH_outW_out = OC * D_out * H_out * W_out
D_outH_outW_out = D_out * H_out * W_out
# Decompose flat index into (batch, channel, depth, height, width)
b_idx = offsets // OC_D_outH_outW_out
residual = offsets % OC_D_outH_outW_out
oc_idx = residual // D_outH_outW_out
spatial_idx = residual % D_outH_outW_out
d_idx = spatial_idx // (H_out * W_out)
hw_idx = spatial_idx % (H_out * W_out)
h_idx = hw_idx // W_out
w_idx = hw_idx % W_out
# Accumulator for output value (per element)
acc = tl.zeros((BLOCK_SIZE,), dtype=tl.float32)
# Loop over all input channels and 3x3x3 kernel positions
for ic in range(IC):
for kd in range(3):
for kh in range(3):
for kw in range(3):
# Compute corresponding input position for each output
d_in = d_idx - kd
h_in = h_idx - kh
w_in = w_idx - kw
# Check input bounds to avoid invalid memory access
in_bounds = (
(d_in >= 0) & (d_in < D_in) &
(h_in >= 0) & (h_in < H_in) &
(w_in >= 0) & (w_in < W_in)
# Compute input tensor offset
input_offsets = (
b_idx * input_batch_stride +
ic * input_channel_stride +
d_in * input_d_stride +
h_in * input_h_stride +
w_in * input_w_stride
# Load input values with masking (zeros for out-of-bounds)
input_val = tl.load(input_ptr + input_offsets, mask=in_bounds, other=0.0)
# Compute weight tensor offset (flipped in transpose)
weight_offsets = (
oc_idx * weight_out_channels_stride +
ic * weight_in_channels_stride +
kd * weight_d_stride +
kh * weight_h_stride +
kw * weight_w_stride
weight_val = tl.load(weight_ptr + weight_offsets)
# Accumulate product into output accumulator
acc += input_val * weight_val
# Store the final output value
output_offsets = (
b_idx * output_batch_stride +
oc_idx * output_channel_stride +
d_idx * output_d_stride +
h_idx * output_h_stride +
w_idx * output_w_stride
# Store with mask for thread-safety
tl.store(output_ptr + output_offsets, acc, mask=mask)
def triton_conv_transpose3d(x: torch.Tensor, weight: torch.Tensor) -> torch.Tensor:
B, IC, D_in, H_in, W_in = x.shape
OC = weight.shape[1]
# Output dimensions for stride=1, padding=0, kernel=3
D_out = D_in + 2
H_out = H_in + 2
W_out = W_in + 2
# Allocate output tensor
output = torch.empty((B, OC, D_out, H_out, W_out), device=x.device, dtype=x.dtype)
# Launch grid size: enough blocks to cover all output elements
BLOCK_SIZE = 256
grid = (triton.cdiv(B * OC * D_out * H_out * W_out, BLOCK_SIZE),)
# Launch the kernel
conv_transpose3d_kernel[grid](
x, weight, output,
B, IC, OC, D_in, H_in, W_in, D_out, H_out, W_out,
x.stride(0), x.stride(1), x.stride(2), x.stride(3), x.stride(4),
weight.stride(0), weight.stride(1), weight.stride(2), weight.stride(3), weight.stride(4),
output.stride(0), output.stride(1), output.stride(2), output.stride(3), output.stride(4),
BLOCK_SIZE
return output
@@
def __init__(
self,
in_channels: int,
out_channels: int,
kernel_size: int,
stride: int = 1,
padding: int = 0,
output_padding: int = 0,
groups: int = 1,
bias: bool = False
super().__init__()
# This implementation is hardcoded for kernel_size=3, stride=1, no padding, no bias
assert (
kernel_size == 3 and stride == 1 and
padding == 0 and output_padding == 0 and
groups == 1 and not bias
self.in_channels = in_channels
self.out_channels = out_channels
# Weight shape: [in_channels, out_channels, 3, 3, 3]
self.weight = $n$.Parameter(torch.empty(in_channels, out_channels, 3, 3, 3))
# Initialize with Kaiming uniform
$n$.init.kaiming_uniform_(self.weight, a=5**0.5)
def forward(self, x: torch.Tensor) -> torch.Tensor:
return triton_conv_transpose3d(x, self.weight)
Problem 35 in KernelBench Level 1: Group Normalization with x1.1 speedup.
@@
def group_norm_reduce_kernel(
x_ptr,
mean_group_ptr,
var_group_ptr,
N, C, H, W, num_groups, M,
BLOCK_SIZE_REDUCE: tl.constexpr
# Program IDs for parallel execution across batch and groups
pid0 = tl.program_id(0) # batch index
pid1 = tl.program_id(1) # group index
group_size = M * H * W
group_start = pid0 * (C * H * W) + pid1 * group_size
# Use float accumulators for numerical stability
sum1 = 0.0
sum2 = 0.0
# Loop over the group in chunks for better cache usage
for i in range(0, group_size, BLOCK_SIZE_REDUCE):
offset = i + tl.arange(0, BLOCK_SIZE_REDUCE)
mask = offset < group_size # avoid out-of-bounds loads
offsets = group_start + offset
# Efficient memory access with masking and zero-padding
chunk = tl.load(x_ptr + offsets, mask=mask, other=0.0)
# Accumulate sum and squared sum
sum1 += tl.sum(chunk, axis=0)
sum2 += tl.sum(chunk * chunk, axis=0)
# Numerically stable variance computation
mean = sum1 / group_size
var = (sum2 - mean * sum1) / group_size
# Store per-group mean and variance
mean_idx = pid0 * num_groups + pid1
tl.store(mean_group_ptr + mean_idx, mean)
tl.store(var_group_ptr + mean_idx, var)
@triton.jit
def group_norm_forward_kernel(
x_ptr,
out_ptr,
mean_group_ptr,
var_group_ptr,
weight_ptr,
bias_ptr,
N, C, H, W, num_groups, M, eps,
BLOCK_SIZE: tl.constexpr
# 1D parallelism across total number of elements
pid = tl.program_id(0)
total_elements = N * C * H * W
block_start = pid * BLOCK_SIZE
offsets = block_start + tl.arange(0, BLOCK_SIZE)
mask = offsets < total_elements # bounds checking
# Flattened indexing to recover n, c from offset
total2 = H * W
total3 = C * total2
n = offsets // total3
rest = offsets - n * total3
c = rest // total2
group_index = c // M # which group the channel belongs to
# Memory load with masking
x_offsets = offsets
x_val = tl.load(x_ptr + x_offsets, mask=mask)
# Load per-sample, per-group mean and variance
mean_idx = n * num_groups + group_index
mean_val = tl.load(mean_group_ptr + mean_idx)
var_val = tl.load(var_group_ptr + mean_idx)
# Load affine transformation parameters per channel
gamma = tl.load(weight_ptr + c)
beta = tl.load(bias_ptr + c)
# Avoid negative variance (numerical safety)
var_val = tl.maximum(var_val, 0)
std = tl.sqrt(var_val + eps)
# Normalize and apply affine transformation
normalized = (x_val - mean_val) / std
out_val = normalized * gamma + beta
# Store result with mask to handle edge threads
tl.store(out_ptr + x_offsets, out_val, mask=mask)
@@
def __init__(self, num_features, num_groups):
@@
self.num_groups = num_groups
self.num_features = num_features
# Learnable affine parameters
self.weight = $n$.Parameter(torch.ones(num_features))
self.bias = $n$.Parameter(torch.zeros(num_features))
def forward(self, x):
x = x.contiguous() # ensure contiguous memory layout for Triton
N, C, H, W = x.shape
M = C // self.num_groups # channels per group
# Allocate buffers for per-group statistics
mean_group = torch.empty((N, self.num_groups), device=x.device)
var_group = torch.empty((N, self.num_groups), device=x.device)
# Launch reduction kernel: one thread per (N, group)
grid_reduce = (N, self.num_groups)
group_norm_reduce_kernel[grid_reduce](
x, mean_group, var_group,
N, C, H, W, self.num_groups, M,
BLOCK_SIZE_REDUCE=1024 # large block size for better throughput
# Allocate output tensor
out = torch.empty_like(x)
total_elements = N * C * H * W
# Launch forward kernel: 1D block across all elements
grid_forward = (triton.cdiv(total_elements, 1024),)
group_norm_forward_kernel[grid_forward](
x, out, mean_group, var_group,
self.weight, self.bias,
N, C, H, W, self.num_groups, M, 1e-5,
BLOCK_SIZE=1024 # tuneable block size
return out
Appendix H Hyperparameters of SFT and RL Training
For data preparation, we use temperature=0.6, top_p=0.95 for DeepSeek-R1 (Guo et al., 2025). To label difficulty, we further label tasks into three difficulty levels using Qwen3-235B-Instruct (Team, 2025) (temperature=0.7, top_p=0.8).
For SFT, we training for 2 epochs with a batch size of 16, a learning rate of $1\times 10^{-5}$, and a maximum sequence length of 12,288 tokens. We train the model for 2 epochs using the VeRL framework (Sheng et al., 2025) with a batch size of 32, a learning rate of $1\times 10^{-6}$, the maximum prompt length $2,048$, and the maximum response length $16,384$. We use 8 NVIDIA A100 80GB GPUs for both SFT and RL training.
Appendix I Additional Experiment Results
I.1 Overall Performance Comparison
Model #Params Level1 Level1 (w/o robust)
valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$ valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$
Qwen3 (base) 88B $73.0$ $40.0$ / $14.0$ $0.0$ / $0.0$ $54.0$ $52.0$ / $15.0$ $0.0$ / $0.0$
Qwen3 1414B $82.0$ $65.0$ / $17.0$ $0.0$ / $0.0$ $66.0$ $71.0$ / $16.0$ $0.0$ / $0.0$
Qwen3 3232B $75.0$ $61.0$ / $16.0$ $2.0$ / $0.0$ $52.0$ $50.0$ / $15.0$ $2.0$ / $0.0$
KernelLLM 88B $42.0$ $40.0$ / $20.0$ $0.0$ / $0.0$ $100.0$ $98.0$ / $29.0$ $2.0$ / $0.0$
AutoTriton 88B $97.0$ $78.0$ / $50.0$ $2.0$ / $1.0$ $100.0$ $95.0$ / $70.0$ $8.0$ / $1.0$
TritonRL (ours) 88B $99.0$ $82.0$ / $56.0$ $5.0$ / $1.0$ $99.0$ $83.0$ / $58.0$ $5.0$ / $1.0$
   w/o RL (SFT only) 88B $97.0$ $88.0$ / $44.0$ $4.0$ / $2.0$ $98.0$ $93.0$ / $47.0$ $5.0$ / $2.0$
GPT-oss 120120B $100.0$ $100.0$ / $74.0$ $7.0$ / $2.0$ $100.0$ $100.0$ / $78.0$ $7.0$ / $2.0$
Claude-3.7 - $99.0$ $99.0$ / $53.0$ $3.0$ / $1.0$ $100.0$ $100.0$ / $64.0$ $8.0$ / $1.0$
DeepSeek-R1 685685B $100.0$ $100.0$ / $66.0$ $6.0$ / $2.0$ $100.0$ $100.0$ / $72.0$ $6.0$ / $2.0$
Table 6: Main results on KernelBench Level 1. All metrics are reported in terms of pass@10 (%). We obtained the best result in model parameter sizes <32<32B. The left side shows the results where the validity of generated codes is verified using the robust verifier, checking both syntax and functionality. The right side (w/o robust) shows the results without the robust verifier, where functionality is not checked.
Model #Params Level2 Level2 (w/o robust)
valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$ valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$
Qwen3 (base) 88B $56.0$ $1.0$ /$0.0$ $0.0$ / $0.0$ $94.0$ $52.0$ / $11.0$ $1.0$ / $0.0$
Qwen3 1414B $35.0$ $24.0$ / $1.0$ $0.0$ / $0.0$ $100.0$ $94.0$ / $65.0$ $14.0$ / $1.0$
Qwen3 3232B $31.0$ $16.0$ / $0.0$ $0.0$ / $0.0$ $90.0$ $73.0$ / $22.0$ $7.0$ / $1.0$
KernelLLM 88B $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $98.0$ $96.0$ / $3.0$ $3.0$ / $1.0$
AutoTriton 88B $70.0$ $3.0$ / $0.0$ $0.0$ / $0.0$ $100.0$ $97.0$ / $76.0$ $15.0$ / $0.0$
TritonRL (ours) 88B $69.0$ $41.0$ / $10.0$ $0.0$ / $0.0$ $100.0$ $88.0$ / $42.0$ $13.0$ / $1.0$
   w/o RL (SFT only) 88B $67.0$ $32.0$ / $6.0$ $0.0$ / $0.0$ $100.0$ $98.0$ / $41.0$ $11.0$ / $1.0$
GPT-oss 120120B $39.0$ $38.0$ / $12.0$ $0.0$ / $0.0$ $100.0$ $99.0$ / $74.0$ $23.0$ / $1.0$
Claude-3.7 - $34.0$ $34.0$ / $12.0$ $1.0$ / $0.0$ $100.0$ $98.0$ / $60.0$ $18.0$ / $1.0$
DeepSeek-R1 685685B $30.0$ $29.0$ / $10.0$ $0.0$ / $0.0$ $100.0$ $98.0$ / $72.0$ $25.0$ / $3.0$
Table 7: Main results on KernelBench level 2 tasks. All metrics are reported in terms of pass@10 (%). We obtained the best result in model parameter sizes <32<32B. The left side shows the results where the validity of generated codes is verified using the robust verifier, checking both syntax and functionality. The right side (w/o robust) shows the results without the robust verifier, where functionality is not checked.
I.2 Pass@k Results
In addition to the pass@10 results shown in the main text, we also report pass@1 and pass@5 results for KernelBench Level 1 and Level 2 tasks in Table 8 and Table 9. These results further demonstrate the strong performance of TritonRL in generating valid, correct, and efficient Triton code across various pass@$k$ metrics.
Model pass@1 pass@5 pass@10
Qwen3 (8B) $24.0$ $14.0$ / $4.0$ $0.0$ / $0.0$ $54.0$ $33.0$ / $14.0$ $0.0$ / $0.0$ $73.0$ $40.0$ / $14.0$ $0.0$ / $0.0$
Qwen3 (14B) $25.0$ $22.0$ / $10.0$ $0.0$ / $0.0$ $66.0$ $53.0$ / $14.0$ $0.0$ / $0.0$ $82.0$ $65.0$ / $17.0$ $0.0$ / $0.0$
Qwen3 (32B) $19.0$ $16.0$ / $4.0$ $0.0$ / $0.0$ $52.0$ $41.0$ / $14.0$ $2.0$ / $0.0$ $75.0$ $61.0$ / $16.0$ $2.0$ / $0.0$
AutoTriton $61.0$ $42.0$ / $18.0$ $0.0$ / $0.0$ $94.0$ $73.0$ / $40.0$ $1.0$ / $1.0$ $97.0$ $78.0$ / $50.0$ $2.0$ / $1.0$
TritonRL (ours) $70.0$ $51.0$ / $21.0$ $0.0$ / $0.0$ $97.0$ $80.0$ / $47.0$ $3.0$ / $1.0$ $99.0$ $82.0$ / $56.0$ $5.0$ / $1.0$
   w/o RL (SFT only) $48.0$ $37.0$ /$12.0$ $0.0$ / $0.0$ $94.0$ $77.0$ / $31.0$ $2.0$ /$1.0$ $97.0$ $88.0$ / $44.0$ $4.0$ / $2.0$
GPT-oss $82.0$ $82.0$ / $38.0$ $3.0$ / $1.0$ $100.0$ $100.0$ / $64.0$ $7.0$ / $2.0$ $100.0$ $100.0$ / $74.0$ $7.0$ / $2.0$
Claude-3.7 $78.0$ $73.0$ / $25.0$ $1.0$ / $1.0$ $99.0$ $98.0$ / $40.0$ $1.0$ / $1.0$ $99.0$ $99.0$ / $53.0$ $3.0$ / $1.0$
DeepSeek-R1 $87.0$ $86.0$ / $29.0$ $1.0$ / $0.0$ $99.0$ $98.0$ / $51.0$ $4.0$ / $2.0$ $100.0$ $100.0$ / $66.0$ $6.0$ / $2.0$
Model pass@1 pass@5 pass@10
Qwen3 (8B) $9.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $30.0$ $1.0$ / $0.0$ $0.0$ / $0.0$ $56.0$ $1.0$ / $0.0$ $0.0$ / $0.0$
Qwen3 (14B) $4.0$ $2.0$ / $1.0$ $0.0$ / $0.0$ $23.0$ $17.0$ / $1.0$ $0.0$ / $0.0$ $35.0$ $24.0$ / $1.0$ $0.0$ / $0.0$
Qwen3 (32B) $4.0$ $3.0$ / $0.0$ $0.0$ / $0.0$ $19.0$ $9.0$ / $0.0$ $0.0$ / $0.0$ $31.0$ $16.0$ / $0.0$ $0.0$ / $0.0$
KernelLLM $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$
AutoTriton $21.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $57.0$ $1.0$ / $0.0$ $0.0$ / $0.0$ $70.0$ $3.0$ / $0.0$ $0.0$ / $0.0$
TritonRL (ours) $16.0$ $10.0$ / $0.0$ $0.0$ / $0.0$ $56.0$ $20.0$ / $4.0$ $0.0$ / $0.0$ $71.0$ $29.0$ / $7.0$ $0.0$ / $0.0$
   w/o RL (SFT only) $15.0$ $8.0$ / $0.0$ $0.0$ / $0.0$ $51.0$ $25.0$ / $5.0$ $0.0$ / $0.0$ $67.0$ $32.0$ / $6.0$ $0.0$ / $0.0$
GPT-oss $9.0$ $8.0$ / $2.0$ $0.0$ / $0.0$ $30.0$ $28.0$ / $7.0$ $0.0$ / $0.0$ $39.0$ $38.0$ / $12.0$ $0.0$ / $0.0$
Claude-3.7 $15.0$ $13.0$ / $1.0$ $0.0$ / $0.0$ $31.0$ $30.0$ / $10.0$ $1.0$ / $0.0$ $34.0$ $34.0$ / $12.0$ $1.0$ / $0.0$
DeepSeek-R1 $6.0$ $4.0$ / $0.0$ $0.0$ / $0.0$ $23.0$ $22.0$ / $4.0$ $0.0$ / $0.0$ $30.0$ $29.0$ / $10.0$ $0.0$ / $0.0$

### v2

@@ G.2 Example of invalid Triton codes in KernelBook
G.3 Example of invalid Triton codes for Level 2 tasks
The following is an example of invalid Triton code generated during RL training of TritonRL. This level 2 task aims to implement a two-branch linear transformation with message passing, structured like a Graph Neural Network (GNN) layer. However, while the code implement linear transformations using a custom Triton kernel, it fails to fully realize the intended GNN functionality with only Triton kernels by using torch.bmm.
@@ G.3 Example of invalid Triton codes for Level 2 tasks
G.4 Invalid Triton code examples
@@
channel_idx = (offsets // stride_channel) %
@@
G.5 SFT and RL Dataset construction with KernelBook
@@
For each given PyTorch reference code, we construct an instruction for a teacher model to generate CoTs and Triton kernels as:
@@
Given the instruction for each PyTorch reference code, we collect (CoT, Triton kernel code) pairs from DeepSeek-R1 and construct outputs for SFT by concatenating the pairs as follows:
@@
Appendix H LLM judge for functional validity verification
To further enhance the functional validity verifier beyond unit tests, we employ an LLM-based judge to identify potential cheating behaviors in the generated Triton code, which may pass unit tests and syntax checks but fail to implement the intended functionality. We utilize Qwen3-235B-Instruct (Team, 2025) as the judge model, prompted to determine if the Triton code genuinely implements the same functionality without relying on high-level PyTorch constructs or hardcoded constants. We design the following prompt that details the functional requirements of Triton kernels for the judge model:
Judge Instruction for Functional Validity Verification Detect whether the following Triton kernel implementation of an operation is cheating or not. A kernel is considered ’cheating’ according to the following rules: 1. **All Core Computation Logic Must Be Written in Triton** - All core computation logic must be implemented in Triton and must not delegate any part of the computation to high-level PyTorch APIs, such as torch.nn or torch.nn.functional. - Wrapping high-level torch APIs in a Triton kernel wrapper without expressing the computation logic in Triton is not acceptable. - Examples: (1). When implementing matrix multiplication in Triton, the use of torch.matmul, @, torch.nn.Linear, or similar high-order APIs is forbidden. (2). When implementing layer normalization in Triton, the use of torch.nn.functional.layer_norm is forbidden. (3). When implementing softmax in Triton, the use of torch.softmax is forbidden. (4). When implementing convolution or convolution transpose, the use of torch.nn.functional.conv2d or torch.nn.ConvTranspose3d is forbidden. (5). When implementing attentions in Triton, the use of transformer library or other high-level APIs is forbidden. 2. **No Trivial Identity or No-Op Computation** - Any part of the Triton kernel should always apply a mathematically-nontrivial operation to the input tensors like computation, reduction, or data layout change. - Examples: (1). Kernels that merely copy input to output (i.e., ’output_val = input_val’) without performing meaningful computation are not allowed. (2). Kernels that does ’output_val = input_val + 0’ or ’* 1.0’ are not allowed. 3. **No Omission of Necessary Computations** - When implementing an operator that involves multiple components, the kernel must implement **all** required parts. Skipping any required mathematical component of the algorithm is considered cheating. - Examples: 1. When implementing LayerNorm, the kernel must include computations of - Mean and variance computation - Normalization - Scaling and bias if applicable 2. When implementing matmul followed by a transpose, the kernel must not omit the transpose of the matmul results. 4. **The Kernel Must Address a Known Performance Bottleneck** - The operation should be **non-trivial** enough that writing a custom kernel could yield meaningful performance benefits using PyTorch API - Examples: - Fused element-wise ops (e.g., GELU + dropout + bias) - Reductions (e.g., softmax, LayerNorm, RMSNorm) - Operations that can benefit from different memory layout of intermediate tensors 5. **The kernel Should Make Efficient Use of Available Parallelism in GPU** - Kernels should make use of block/thread-level parallelism (’program_id’, ’arange’, ’BLOCK_SIZE’, etc.) - No scalar-only computations or logic that doesn’t scale with input size
Appendix I Evaluation and examples
I.1 Evaluation with KernelBench
@@
I.2 Example of Triton codes generated by TritonRL (D) with speedup >1>1
@@
Problem 94 in KernelBench Level 1: MSE (Mean Squared Error) with x3 speedup.
@@
def mse_kernel(
pred_ptr,
target_ptr,
output_ptr,
n_elements,
@@
# 1. Block Setup
pid = tl.program_id(axis=0)
@@
# 2. Masked Load with Zero-Padding
# CRITICAL: other=0.0 ensures that elements outside the valid range
# contribute 0 to the sum, preventing calculation errors.
pred_vals = tl.load(pred_ptr + offsets, mask=mask, other=0.0)
target_vals = tl.load(target_ptr + offsets, mask=mask, other=0.0)
# 3. Compute Squared Error
# This happens in parallel across all threads in the block
diff = pred_vals - target_vals
sq_diff = diff * diff
# 4. Intra-Block Reduction
# tl.sum collapses the 1024 values in this block into a single scalar.
# This is extremely fast as it uses shared memory/registers, avoiding VRAM.
block_sum = tl.sum(sq_diff, axis=0)
# 5. Global Accumulation
# We use an atomic add to aggregate the partial sums from each block
# into the single global output location.
tl.atomic_add(output_ptr, block_sum)
def triton_mse(pred: torch.Tensor, target: torch.Tensor) -> torch.Tensor:
# Ensure data layout is contiguous for coalesced memory access
if not pred.is_contiguous(): pred = pred.contiguous()
if not target.is_contiguous(): target = target.contiguous()
n_elements = pred.numel()
if n_elements == 0:
return torch.tensor(0.0, device=pred.device)
# Output must be initialized to zero because atomic_add accumulates on top of it
output = torch.zeros(1, device=pred.device, dtype=pred.dtype)
BLOCK_SIZE = 1024
# Standard grid calculation
grid = lambda meta: (triton.cdiv(n_elements, meta[’BLOCK_SIZE’]),)
mse_kernel[grid](pred, target, output, n_elements, BLOCK_SIZE=BLOCK_SIZE)
# Final division by N happens on Python side (negligible cost)
return output[0] / n_elements
@@
def forward(self, predictions, targets):
return triton_mse(predictions, targets)
Problem 99 in KernelBench Level 1: Triplet Margin Loss with x4 speedup.
@@
def triplet_margin_loss_kernel(
anchor_ptr,
positive_ptr,
negative_ptr,
output_ptr,
margin,
# FEATURE_SIZE as constexpr allows the compiler to optimize the loop unrolling
FEATURE_SIZE: tl.constexpr,
BLOCK_SIZE: tl.constexpr,
# 1. Parallelize by Batch (Row)
# Each program instance (block) handles exactly one triplet sample from the batch.
pid = tl.program_id(axis=0)
# Calculate starting memory addresses for this specific row
offset_start = pid * FEATURE_SIZE
anchor_row = anchor_ptr + offset_start
positive_row = positive_ptr + offset_start
negative_row = negative_ptr + offset_start
# Registers to accumulate squared distances across the loop
ap_sum = 0.0
an_sum = 0.0
# 2. Tiled Processing of Features
# We loop over the feature dimension in chunks (tiles) of BLOCK_SIZE.
# This keeps register pressure low and allows handling features > 1024 dims.
for j in range(0, FEATURE_SIZE, BLOCK_SIZE):
# Generate offsets for the current tile
offsets = j + tl.arange(0, BLOCK_SIZE)
mask = offsets < FEATURE_SIZE
# Load data chunks (coalesced reads)
a = tl.load(anchor_row + offsets, mask=mask, other=0.0)
p = tl.load(positive_row + offsets, mask=mask, other=0.0)
n = tl.load(negative_row + offsets, mask=mask, other=0.0)
# Vectorized math on the chunk
diff_ap = a - p
sq_ap = diff_ap * diff_ap
diff_an = a - n
sq_an = diff_an * diff_an
# 3. Partial Reduction
# Reduce the chunk to a scalar and add to the running total.
# tl.sum is highly optimized for intra-block communication.
ap_sum += tl.sum(sq_ap, axis=0)
an_sum += tl.sum(sq_an, axis=0)
# 4. Final Loss Calculation (per row)
d_ap = tl.sqrt(ap_sum)
d_an = tl.sqrt(an_sum)
loss_i = d_ap - d_an + margin
loss_i = tl.where(loss_i > 0, loss_i, 0.0) # ReLU
# 5. Global Accumulation
# Aggregate this row’s loss into the total batch loss.
# Note: For massive batch sizes, this atomic_add can become a contention hotspot.
tl.atomic_add(output_ptr, loss_i)
@@
def __init__(self, margin=1.0):
@@
self.margin = margin
def forward(self, anchor, positive, negative):
# Ensure contiguous memory for safe pointer arithmetic
anchor = anchor.contiguous()
positive = positive.contiguous()
negative = negative.contiguous()
batch_size, n_features = anchor.shape
# Allocate output tensor (scalar)
total_loss = torch.zeros(1, device=anchor.device, dtype=anchor.dtype)
# Grid Size = Batch Size (one block per sample)
grid = (batch_size,)
triplet_margin_loss_kernel[grid](
anchor,
positive,
negative,
total_loss,
self.margin,
FEATURE_SIZE=n_features, # Pass actual dimension dynamically
BLOCK_SIZE=1024, # 1024 is standard for max occupancy
return total_loss[0] / batch_size